# The impact of non-Gaussianity when searching for Primordial Black Holes with LISA

### v1

@@ 1 Introduction
Among other sources, scalar-induced gravitational waves (SIGW) [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] constitute a particularly interesting cosmological GW background detectable by LISA (see also [27] for a recent physical explanation of the origin of the signal). As the name suggests, these signals are generated from scalar fluctuations, which act as sources for tensor modes at second order in perturbation theory. As the shape and amplitude of the resulting SIGW spectrum are derived from the scalar power spectrum [28], its detection would be fundamental to unveil the dynamics of the early-Universe on scales much smaller than the ones probed by Planck and more recently by the ACT and SPT collaborations [29, 30, 31]. Various inflationary models [32, 33, 34] [^1] Examples of this are single-field scenarios where the potential shows specific features which lead to an ultra-slow-roll (USR) phase [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] as well as multi-field models [32, 33, 46, 47, 48, 49, 50, 51, 52, 53], but also mechanisms such as preheating or early matter-dominated eras which can produce a strong enhancement of the GW spectrum without the need of a largely enhance of the scalar power spectrum [54, 55, 56, 57, 58]., in fact, predict an amplification of the scalar fluctuations around LISA scales with respect to the nearly-flat power spectrum on CMB scales [59] and this enhancement could possibly generate a large amount of SIGW. In addition, SIGWs are directly sourced by the four-point function of the scalar fluctuation and are sensitive to the possible presence of non-Gaussianity (NG) in the scalar perturbation, as it has been recently analysed by [60, 23, 24, 61]. Primordial NG represents a further powerful probe of the early Universe, being sensitive to the dynamics and interactions in the early Universe. Although standard inflationary scenarios predict mild deviations from Gaussian behaviour, other models could generate substantial NG perturbations [62, 63, 64]; thus, constraining primordial NG is fundamental for discriminating between models. Nevertheless, Planck has been able to put tight constraints on the amount of NG, typically parametrized by the quantity $f_{\rm NL}$, on large scales [65], but small scales remain unconstrained.
@@ 2.2 Perturbativity
with $F_{\rm NL,1}=1$ and $F_{\rm NL,2}=(3/5)f_{\rm NL}$, etc. The coefficients (2.17) can then be expressed by the series
@@ 2.3 Primordial black hole abundance
Looking at the amplitude required to achieve $f_{\rm PBH}\simeq 1$, we observe that the maximum achievable amplitude $A$ of the power spectrum typically peaks around $f_{\mathrm{NL}}\approx-2$, in agreement with Ref. [121, 122], with only a mild sensitivity to the spectral width, especially for positive NG. Notably, the curves in Fig. 2 display a steep slope between the purely Gaussian case and moderately positive NG (up to $f_{\mathrm{NL}}\approx 10$). This implies that even relatively small deviations from Gaussianity can have a significant effect on the resulting PBH abundance. This is due to the exponential sensitivity of PBH formation to the tail of the probability distribution for curvature perturbations. As a result, mild positive NG can amplify the abundance by several orders of magnitude, an effect that is especially pronounced for narrow power spectra, where the contribution from rare, large fluctuations is dominant. This sensitivity underscores the importance of accurately modelling non-Gaussian effects when translating constraints from GW observations into bounds on PBH dark matter; particularly, as we will see in the next sections, in the asteroid-mass window probed by experiments such as LISA. Moreover we stress that adopting the local expansion to approximate the imprints of NG for some inflation models leads to a large error of many orders of magnitude in the PBH abundance estimate [81].
@@ 4.1 Sensitivity to PBH abundance
We begin by highlighting that the evaporation constraints for extended mass functions differ significantly from those in the monochromatic case, due to the finite width of the mass distribution [117]. While monochromatic scenarios typically impose evaporation bounds up to $\langle M_{\rm PBH}\rangle\simeq 10^{-16}$ $M_{\odot}$ (see also [149, 150, 151, 152]; and constraints from EDGES [153], CMB [154], INTEGRAL [155, 156], the 511 keV line [157, 158], Voyager [159] and EGRB [160]), the presence of a critical tail in extended mass functions pushes these limits further, up to $\langle M_{\rm PBH}\rangle\simeq 4\times 10^{-16}M_{\odot}$ as can be seen in Fig. 6. This shift significantly affects the minimum level of primordial NGs required to preserve any viable parameter space in which the entirety of dark matter could consist of PBHs. The constraints for heavier PBHs, which are dominated by microlensing, including from Hyper Suprime-Cam [161] and OGLE [162, 163, 164], do exclude slightly lower masses when the width of the mass function has been accounted for, although the effect is milder.
@@
Such prospective constraints can be relaxed when $f_{\rm NL}$ sufficiently large, allowing PBHs with masses $\sim 10^{-7}M_{\odot}$ to remain viable in the case of non-detection of a SIGW by LISA. This mass range could be probed by future interferometers such as the Einstein Telescope (ET) [167, 5, 168].
@@ 5 Conclusions
LISA’s sensitivity to the power spectrum translates into prospective constraints on the PBH abundance. In the Gaussian case, a non-detection of SIGWs in the LISA band would essentially exclude any conceivable amount of PBHs in the asteroid-mass range ($10^{-17}M_{\odot}\lesssim M_{\mathrm{PBH}}\lesssim 10^{-5}M_{\odot}$). With sizable positive non-Gaussianity ( $f_{\mathrm{NL}}>5$), the limits loosen substantially, leaving room for PBHs to comprise all of dark matter in the low mass end of the asteroid mass window. Astrophysical foregrounds can also soften the prospective constraints, but their effect is much milder when compared to the effect of non-Gaussianity.
@@ Acknowledgments
We thank A. Ricciardone, A. Riotto, and V. Vaskonen for useful discussions. H.V. and G.P. are supported by the Estonian Research Council grants PSG869, RVTT7, KOHTO34 and the Center of Excellence program TK202. G.P. acknowledges partial financial support by ASI Grant No. 2016-24-H.0 and thanks Fondazione Angelo Della Riccia and Fondazione Aldo Gini for financial support.

### v2

@@ 1 Introduction
Among other sources, scalar-induced gravitational waves (SIGW) [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] constitute a particularly interesting cosmological GW background detectable by LISA (see also [27] for a recent physical explanation of the origin of the signal). As the name suggests, these signals are generated from scalar fluctuations, which act as sources for tensor modes at second order in perturbation theory. As the shape and amplitude of the resulting SIGW spectrum are derived from the scalar power spectrum [28], its detection would be fundamental to unveil the dynamics of the early-Universe on scales much smaller than the ones probed by Planck and more recently by the ACT and SPT collaborations [29, 30, 31]. Various inflationary models [32, 33, 34] [^1] Examples of this are single-field scenarios where the potential shows specific features which lead to an ultra-slow-roll (USR) phase [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] as well as multi-field models [32, 33, 46, 47, 48, 49, 50, 51, 52, 53], but also mechanisms such as preheating or early matter-dominated eras which can produce a strong enhancement of the GW spectrum without the need of a largely enhance of the scalar power spectrum [54, 55, 56, 57, 58]., in fact, predict an amplification of the scalar fluctuations around LISA scales with respect to the nearly-flat power spectrum on CMB scales [59] and this enhancement could possibly generate a large amount of SIGW. In addition, SIGWs are directly sourced by the four-point function of the scalar fluctuation and are sensitive to the possible presence of non-Gaussianity (NG) in the scalar perturbation, as it has been recently analysed by [60, 61, 23, 24, 62]. Primordial NG represents a further powerful probe of the early Universe, being sensitive to the dynamics and interactions in the early Universe. Although standard inflationary scenarios predict mild deviations from Gaussian behaviour, other models could generate substantial NG perturbations [63, 64, 65]; thus, constraining primordial NG is fundamental for discriminating between models. Nevertheless, Planck has been able to put tight constraints on the amount of NG, typically parametrized by the quantity $f_{\rm NL}$, on large scales [66], but small scales remain unconstrained.
@@ 2.2 Perturbativity
with $F_{\rm NL,1}=1$ and $F_{\rm NL,2}=(3/5)f_{\rm NL}$, etc. The coefficients (2.17) can then be expressed by the series by analytically evaluating the integrals (see [106] for an alternative derivation)
@@ 2.3 Primordial black hole abundance
Looking at the amplitude required to achieve $f_{\rm PBH}\simeq 1$, we observe that the maximum achievable amplitude $A$ of the power spectrum typically peaks around $f_{\mathrm{NL}}\approx-2$, in agreement with Ref. [123, 124], with only a mild sensitivity to the spectral width, especially for positive NG. Notably, the curves in Fig. 2 display a steep slope between the purely Gaussian case and moderately positive NG (up to $f_{\mathrm{NL}}\approx 10$). This implies that even relatively small deviations from Gaussianity can have a significant effect on the resulting PBH abundance. This is due to the exponential sensitivity of PBH formation to the tail of the probability distribution for curvature perturbations. As a result, mild positive NG can amplify the abundance by several orders of magnitude, an effect that is especially pronounced for narrow power spectra, where the contribution from rare, large fluctuations is dominant. This sensitivity underscores the importance of accurately modelling non-Gaussian effects when translating constraints from GW observations into bounds on PBH dark matter; particularly, as we will see in the next sections, in the asteroid-mass window probed by experiments such as LISA. Moreover we stress that adopting the local expansion to approximate the imprints of NG for some inflation models leads to a large error of many orders of magnitude in the PBH abundance estimate [82, 125, 126].
@@ 4.1 Sensitivity to PBH abundance
[…]
@@
Such prospective constraints can be relaxed when $f_{\rm NL}$ sufficiently large, allowing PBHs up to masses $\sim 10^{-9}M_{\odot}$ ($\sim 10^{-10}M_{\odot}$) with $f_{\rm NL}=5$ (50) to remain viable in the case of non-detection of a SIGW by LISA. In absence of NGs, the constraints are limited at masses $\sim 10^{-7}M_{\odot}$. A mass range that could be probed by future interferometers such as the Einstein Telescope (ET) [172, 5, 173].
@@ 5 Conclusions
LISA’s sensitivity to the power spectrum translates into prospective constraints on the PBH abundance. In the Gaussian case, a non-detection of SIGWs in the LISA band would essentially exclude any conceivable amount of PBHs in the asteroid-mass range ($10^{-16}M_{\odot}\lesssim M_{\mathrm{PBH}}\lesssim 10^{-7}M_{\odot}$). With sizable positive non-Gaussianity ( $f_{\mathrm{NL}}\gtrsim 5$), the limits loosen substantially, leaving room for PBHs to comprise all of dark matter in the low mass end of the asteroid mass window. Astrophysical foregrounds can also soften the prospective constraints, but their effect is much milder when compared to the effect of non-Gaussianity.
@@ Acknowledgments
We thank A. Ricciardone, A. Riotto, and V. Vaskonen for useful discussions. H.V. and G.P. are supported by the Estonian Research Council grants PSG869, RVTT7, KOHTO34, TARISTU24-TK3, TARISTU24-TK10 and the Center of Excellence program TK202. G.P. acknowledges partial financial support by ASI Grant No. 2016-24-H.0 and thanks Fondazione Angelo Della Riccia and Fondazione Aldo Gini for financial support.